OpenStudy (anonymous):

Show that the function y=Ae^-x + Be^-x satisfies the differential equation y"+2y'+y=0

OpenStudy (aakashsudhakar):

y = A(e^-x) + B(e^-x) y' = -A(e^-x) - B(e^-x) = -1[A(e^-x) + B(e^-x)] = -y y" = A(e^-x) + B(e^-x) = +y y = +y y' = -y y" = +y y" + 2y' + y = 0 (+y) + 2(-y) + (+y) = 0 y - 2y + y = 0 2y - 2y = 0 0y = 0 0 = 0 Or, the long way... y" + 2y' + y = 0 (A(e^-x) + B(e^-x)) + 2(-A(e^-x) - B(e^-x)) + (A(e^-x) + B(e^-x)) = 0 2A(e^-x) + 2B(e^-x) - 2A(e^-x) - 2B(e^-x) = 0 0A(e^-x) + 0B(e^-x) = 0 0 = 0 Therefore, y = A(e^-x) + B(e^-x) does satisfy the differential equation y" + 2y' + y = 0.